# Adaptive RKHS Fourier Features for Compositional Gaussian Process Models

### v1

@@ 5 Experiments
One of our goals is to evaluate how the performance of IDDGPs and DLFMs varies with the number of global Fourier features and the effect of the ODE parameters $\alpha$ and $\beta$ on the learning process. Fig. 4(a) illustrates the progressions of performance metrics, e.g., test Root Mean Square Error (RMSE) and Negative Marginal Log-Likelihood (NMLL) using 100 and 400 inducing frequencies. The yellow lines of DLFMs align closely with the IDDGPs’ blue line during the first 2000 iterations, where DLFMs maintain a fixed small $\beta$ value ($\beta=10^{-6},\alpha=1$). We use this setting to illustrate how our DLFM with VFRFs can replicate the original IDDGPs as expected. After $2000$ iterations, we allow optimisation of $\alpha$ and $\beta$, leading to subsequent improvements in the testing metrics, and suggesting continuous learning with ODE-based Fourier features. Additionally, optimising all parameters from the beginning (red lines) yields the best results across various setups. Fig. 4(b) compares the performance of different models with all parameters optimised from the beginning. It is unsurprising that increasing the number of inducing frequencies typically results in better performance. The results reveal that while RFF-based DLFMs exhibit the lowest NMLL, they show the highest RMSE, reflecting a lack of precision. DLFMs equipped with VFRFs consistently surpass both the DGP with local inducing points and the IDDGPs in both RMSE and NMLL, highlighting our model’s enhanced ability to accurately capture the global structure and non-stationarity of the data.
UCI Regression Benchmarks
To demonstrate the versatility and effectiveness of our model on domain-agnostic real-world data, we conduct evaluations on five diverse regression datasets from the UCI Machine Learning Repository (Dua and Graff, 2019). These datasets vary in size and feature dimensionality, allowing us to test the model’s adaptability across different scenarios (see Fig. 5). Consistent with standard practice (Salimbeni and Deisenroth, 2017), our regression tasks involve multivariate inputs and a univariate target. We reserve 10% of each dataset for testing, normalise the inputs to the range $[0,3]$, and standardise the outputs based on the mean and standard deviation of the training set (these transformations are then reversed for evaluation). All models run with Matérn-$\frac{3}{2}$ kernels, employing 20 inducing points or frequencies and three output dimensions per layer. We maintained the same hyperparameters and initialisation across all tests. The figure illustrates that our models achieve comparable performance to the baselines. Notably, our models with two layers outperform IDDGP counterparts on the Energy, Power, and Naval datasets. We also observe that increasing the number of layers generally enhances the model’s representational capacity, resulting in improved performance.
@@ 6 Conclusion
In this work, we proposed a variational approach to incorporating RKHS Fourier features in compositional GP models. Particularly, we designed a set of adaptive inter-domain Fourier features for hierarchical LFMs with modified ODEs. Our empirical evaluations across various datasets demonstrate that our model extends inter-domain GPs with RKHS Fourier features and has an enhanced ability compared to conventional GPs with local approximation in fitting complex data with non-stationary and global patterns. Future work will focus on extending our model’s use to other challenging machine-learning tasks requiring the integration of specific domain knowledge of higher-order ODEs.
@@ 7 Model Comparison
$\mathbf{Feature}$$\mathbf{Physics}$-$\mathbf{informed}$$\mathbf{Structure}$$Inducing\ Points$$RFF$$VFF$$no$$yes$$shallow$$deep$SVGP [Titsias, 2009]LFM [Alvarez et al., 2009]DGP [Damianou and Lawrence, 2013][Salimbeni and Deisenroth, 2017]SSGP[Lázaro-Gredilla et al., 2010][Gal and Turner, 2015]DGP-RFF[Cutajar et al., 2017]DLFM-RFF[McDonald and Álvarez, 2021]LFM-RFF[Guarnizo and Álvarez, 2018]IDDGP[Rudner et al., 2020]GP-VFF[Hensman et al., 2018]This work
@@
8 Variational Fourier Feature
Variational Fourier Features (VFFs, [15]) are designed on a Matérn Reproducing Kernel Hilbert Space (RKHS). Specifically, Matérn-$\frac{1}{2}/\frac{3}{2}/\frac{5}{2}$ kernels with inputs $x,x^{\prime}\in\mathbb{R}$ are of particular interest:
@@ 8 Variational Fourier Feature
where $r=|x-x^{\prime}|$, $\sigma^{2}$ is the kernel’s output-scale (or variance), and $l$ is the length-scale. We reiterate the closed-form RKHS inner products for Matérn-$\frac{1}{2}$ on $[a,b]$ here ( for other Matérn kernels and more details see Durrande et al. [7]) :
@@
but also make it feasible to complete the cross-covariance if $t$ is outside $[a,b]$. In Appx. 10, we further utilise the conclusion of VFFs in [15] for $t\in\mathbb{R}$ to calculate our Variational Fourier Response Features (VFRFs).
@@
9 LFMs for First-Order Dynamical System
@@
where a convolutional operator ${G}$ acting on $u$ is represented as $f=G\circ u$. Conventional LFMs establish a GP over $f\sim\mathcal{GP}(0,G\circ k\circ G)$, where $k\circ G$ signifies $G$ operating the second argument of the kernel. The LFM kernels are computed analytically in [17, 1], but their expressions are based on RBF kernels. In the subsequent part, we will present the closed-form covariance expressions $G\circ k\circ G$ for the Matérn-$\frac{1}{2}/\frac{3}{2}/\frac{5}{2}$ kernels, respectively. All analytical LFM covariance functions discussed in this work are summarised in Table 1 and illustrated in the left panel of Fig. 1, 7 and 8. Furthermore, we introduce the approximation of our LFM kernels using random Fourier features in Appx. 9.2.
9.1 Analytical LFM Matérn  Kernels
@@
9.2 Random Fourier Approximation of Kernels
Stationary kernels can be approximated using random Fourier features [24] using Bochner’s theorem [27]. In the illustrative experiment, we approximate the Matérn LFM covariance proposed in our work using random Fourier features (also termed Random Fourier Response Features (RFRFs) by [11]). The random Fourier features of the LFM with a Matérn-$\nu\ (\nu=\tfrac{1}{2}/\tfrac{3}{2}/\tfrac{5}{2})$ kernel is given by
@@ 9.2 Random Fourier Approximation of Kernels
10 Variational Fourier Response Features for LFMs
@@
We detail the derivation of Matérn-$\frac{1}{2}$ (listed in Table 3 and 4) and directly give the results of Matérn-$\frac{3}{2}$ (see Table 6 and 7) and Matérn-$\frac{5}{2}$ (see Table 9 and 10). These expressions of VFRFs will revert to VFFs under certain conditions. The LFM kernel and the VFRFs for Matérn-$\frac{3}{2}$ and Matérn-$\frac{5}{2}$ with the same hyperparameters in the main text are depicted in Fig. 7 and Fig. 8, respectively.
@@
10.1 Matérn-$\frac{1}{2}$ Cross-covariance
@@
The location of $\tau$ determines the expression of $h(z,\tau^{\prime})=\langle\phi(\cdot;z),k(\cdot,\tau^{\prime})\rangle_{\mathcal{H}}$ (i.e., $P\circ k$) [15], as summarized in the subsequent table (For Matérn-$\frac{3}{2}/\frac{5}{2}$, see Table 5/8). The columns of the tables indicate the VFFs for input $\tau^{\prime}$ located inside/outside $[a,b]$. The cross-covariance can be derived by substituting $h(z,\tau^{\prime})$ in (33). Table 3 and Table 4 collect the VFRFs for Matérn-$\frac{1}{2}$ LFMs with cosine/sine projection basis functions at different locations of $t^{\prime}$ in (33).
@@ 10.1.2 Sine Features ($\phi(t;z)=\sin(z(t-a))$)
The VFRFs of the LFM with a Matérn-$\frac{1}{2}$ latent force are summarised in Table 3 and 4, where the absolute distances to the interval ends are denoted as $r_{a}=|t^{\prime}-a|$ and $r_{b}=|t^{\prime}-b|$ and the phase shift is $\theta=-\arctan(\frac{z}{\gamma})$. The features are continuous at $\gamma=\lambda$ when $t^{\prime}>b$.
@@
10.2 Matérn-$\frac{3}{2}/\frac{5}{2}$ Cross-covariance
Based on Table 5 and 8, we give the VFRFs for Matérn-$\frac{3}{2}$ and Matérn-$\frac{5}{2}$ LFMs with $\theta=-\arctan(\frac{z}{\gamma})$ in Table 6,7,9 and 10. Also, the derived cross-covariance expressions can return to the scaled VFFs of the latent force.
@@ 10.2 Matérn-$\frac{3}{2}/\frac{5}{2}$ Cross-covariance
11 Experimental Details
All models in the experimental section are implemented using GPyTorch [10], trained by Adam Optimizer on an NVIDIA A100-SXM4 GPU (for TIMIT and UCI datasets) or Apple Macbook CPUs (for illustrative examples), with a learning rate of 0.01 and a batch size of 10,000. The models using doubly stochastic variational inference, e.g., IDDGPs, DLFM-VFRF, employ five samples for layer-wise sampling during training. We follow [29] to set up a linear mean function for all the inner layers and a zero-mean function for the outer layer to avoid pathological behaviour [8]. The weights of the linear mean function are fixed and determined by SVD if the input and output dimensions are not equal. The variational distributions over inducing variables are initialised to normal distributions with zero mean and variances identity for the outer layers and $10^{-5}$ for the inner layers. The inducing points are initialised with K-means. All models, including RFF-based models, used 100 Monte Carlo samples for evaluations on test data.

### v2

@@ 5.2 TIMIT Speech Signals
One of our goals is to evaluate how the performance of IDDGPs and DLFMs varies with the number of global Fourier features and the effect of the ODE parameters $\alpha$ and $\beta$ on the learning process. Fig. 5(a) illustrates the progressions of performance metrics, e.g., test RMSE and NMLL using 100 and 400 inducing frequencies. The yellow lines of DLFMs align closely with the IDDGP’s blue line during the first 2000 iterations, where DLFMs maintain a fixed small $\beta$ value ($\beta=10^{-6},\alpha=1$). We use this setting to illustrate how our DLFM-VFRF can replicate the original IDDGPs as expected when $\beta\rightarrow 0^{+}$. After $2000$ iterations, we allow optimisation of $\alpha$ and $\beta$, leading to subsequent improvements in the testing metrics, and suggesting continuous learning with ODE-based Fourier features. Additionally, optimising all parameters from the beginning (red lines) yields the best results across various setups. Fig. 5(b) compares the performance of different models with all parameters optimised from the beginning. It is unsurprising that increasing the number of inducing frequencies typically results in better performance. The results reveal that while RFF-based DLFMs exhibit the lowest NMLL, they show the highest RMSE, reflecting a lack of precision on test data points. DLFMs equipped with VFRFs consistently surpass both the DGP with local inducing points and the IDDGPs in terms of both RMSE and NMLL, highlighting our model’s enhanced ability to accurately capture the global structure and non-stationarity of the data.
Running Time Comparison
Theoretically, the extra running time of our model compared to IDDGPs mainly lies in the more complex forward computation on covariance entries and the backward gradient update on the extra ODE parameters. We record the wall-clock running time (per iteration) of models with the number of inducing frequencies ranging from 100 to 400 in Table 2 below. The results are averaged over five runs, and we exclude the standard deviations as they are quite small.
Table 2: Wall-clock training time of IDDGPs and our DLFM-VFRFs with different numbers of inducing frequencies $M$.
Model/$M$ 100 200 300 400
IDDGP-$M$ 0.279s 0.633s 1.138s 1.891s
Ours-$M$ 0.313s 0.686s 1.207s 1.977s
From Table 2, we observe that our model incurs slightly higher runtime overhead compared to IDDGP. Despite this, the runtime difference compared with IDDGP with VFFs remains acceptable even with 400 inducing features, especially given the improved flexibility and modelling capacity of our approach. The runtime gap might be reduced with some computational optimization techniques (e.g., JIT) implemented.
5.3 UCI Regression Benchmarks
@@
To demonstrate the versatility and effectiveness of our model on domain-agnostic real-world data, we conduct evaluations on five diverse regression datasets from the UCI Machine Learning Repository (Dua and Graff,, 2019). These datasets vary in size and feature dimensionality, allowing us to test the model’s adaptability across different scenarios (see Fig. 6). Consistent with standard practice (Salimbeni and Deisenroth,, 2017), our regression tasks involve multivariate inputs and a univariate target. We reserve 10% of each dataset for testing, normalise the inputs to the range $[0,3]$, and standardise the outputs based on the mean and standard deviation of the training set (these transformations are then reversed for evaluation). Following the setups in Rudner et al., (2020) and McDonald and Álvarez, (2021), all models run with Matérn-$\frac{3}{2}$ kernels and employ 20 inducing points or frequencies. We employ three output dimensions per layer. We maintained the same experimental settings and initialisation across all tests. The figure illustrates that our models achieve comparable performance to the baselines. Notably, our models with two layers outperform IDDGP counterparts on the Energy, Power, and Naval datasets. We also observe that increasing the number of layers generally enhances the model’s representational capacity, resulting in improved performance.
@@ 6 CONCLUSION
In this work, we adapt VFFs to the latent force framework, which inherently involves convolution operators with Green’s functions. This adaptation introduces flexibility in modelling dynamics while preserving computational traceability. By introducing trainable parameters in the Green’s function, we provide a mechanism for dynamically adjusting the inter-domain features. We further employ the inter-domain Fourier features in hierarchical LFMs. Our empirical evaluations across various datasets demonstrate that our model extends inter-domain GPs with RKHS Fourier features and has enhanced their modelling capacity for non-stationary and global structures.
Limitations and Future Work
The current experiments are only based on models from first-order ODEs. Besides, computing the piece-wise VFRFs at intermediate layers may result in extra computational costs. Future work will focus on developing a normalization method at intermediate layers to accelerate inference and on extending our model’s use to other challenging machine-learning tasks requiring the integration of specific domain knowledge of higher-order ODEs. Extending DLFMs to incorporate other recently proposed Fourier features, such as those in Cheema and Rasmussen, (2024), represents a promising direction.
Acknowledgements
We are grateful for the computational resources from UoM CSF3, UK, and appreciate the insightful discussions with Xiaoyu Jiang. X.S. was supported by the UoM-CSC Joint Scholarship.
@@ Appendix A Model Comparison
$\mathbf{Feature}$$\mathbf{Physics}$-$\mathbf{informed}$$\mathbf{Structure}$$Inducing\ Points$$RFF$$VFF$$no$$yes$$shallow$$deep$SVGP (Titsias,, 2009)LFM (Alvarez et al.,, 2009)DGP (Damianou and Lawrence,, 2013)(Salimbeni and Deisenroth,, 2017)SSGP(Lázaro-Gredilla et al.,, 2010)(Gal and Turner,, 2015)DGP-RFF(Cutajar et al.,, 2017)DLFM-RFF(McDonald and Álvarez,, 2021)LFM-RFF(Guarnizo and Álvarez,, 2018)IDDGP(Rudner et al.,, 2020)GP-VFF(Hensman et al.,, 2018)This work
@@
Appendix B Variational Fourier Feature
Variational Fourier Features (VFFs, (Hensman et al.,, 2018)) are designed on a Matérn Reproducing Kernel Hilbert Space (RKHS). Specifically, Matérn-$\frac{1}{2}/\frac{3}{2}/\frac{5}{2}$ kernels with inputs $x,x^{\prime}\in\mathbb{R}$ are of particular interest:
@@ Appendix B Variational Fourier Feature
where $r=|x-x^{\prime}|$, $\sigma^{2}$ is the kernel’s output-scale (or variance), and $l$ is the length-scale. We reiterate the closed-form RKHS inner products for Matérn-$\frac{1}{2}$ on $[a,b]$ here ( for other Matérn kernels and more details see Durrande et al., (2016)) :
@@
but also make it feasible to complete the cross-covariance if $t$ is outside $[a,b]$. In Appx. D, we further utilise the conclusion of VFFs in Hensman et al., (2018) for $t\in\mathbb{R}$ to calculate our Variational Fourier Response Features (VFRFs).
@@
Appendix C LFMs for First-Order Dynamical System
@@
where a convolutional operator ${G}$ acting on $u$ is represented as $f=G\circ u$. Conventional LFMs establish a GP over $f\sim\mathcal{GP}(0,G\circ k\circ G)$, where $k\circ G$ signifies $G$ operating the second argument of the kernel. The LFM kernels are computed analytically in Lawrence et al., (2006); Alvarez et al., (2009), but their expressions are based on RBF kernels. In the subsequent part, we will present the closed-form covariance expressions $G\circ k\circ G$ for the Matérn-$\frac{1}{2}/\frac{3}{2}/\frac{5}{2}$ kernels, respectively. All analytical LFM covariance functions discussed in this work are summarised in Table 3 and illustrated in the left panel of Fig. 1, 8 and 9. Furthermore, we introduce the approximation of our LFM kernels using random Fourier features in Appx. C.2.
C.1 Analytical LFM Matérn  Kernels
@@
C.2 Random Fourier Approximation of Kernels
Stationary kernels can be approximated using random Fourier features (Rahimi and Recht,, 2007) using Bochner’s theorem (Rudin,, 2017). In the illustrative experiment, we approximate the Matérn LFM covariance proposed in our work using random Fourier features (also termed Random Fourier Response Features (RFRFs) by (Guarnizo and Álvarez,, 2018)). The random Fourier features of the LFM with a Matérn-$\nu\ (\nu=\tfrac{1}{2}/\tfrac{3}{2}/\tfrac{5}{2})$ kernel is given by
@@ C.2 Random Fourier Approximation of Kernels
Appendix D Variational Fourier Response Features for LFMs
@@
We detail the derivation of Matérn-$\frac{1}{2}$ (listed in Table 5 and 6) and directly give the results of Matérn-$\frac{3}{2}$ (see Table 8 and 9) and Matérn-$\frac{5}{2}$ (see Table 11 and 12). These expressions of VFRFs will revert to VFFs under certain conditions. The LFM kernel and the VFRFs for Matérn-$\frac{3}{2}$ and Matérn-$\frac{5}{2}$ with the same hyperparameters in the main text are depicted in Fig. 8 and Fig. 9, respectively.
@@
D.1 Matérn-$\frac{1}{2}$ Cross-covariance
@@
The location of $\tau$ determines the expression of $h(z,\tau^{\prime})=\langle\phi(\cdot;z),k(\cdot,\tau^{\prime})\rangle_{\mathcal{H}}$ (i.e., $P\circ k$) (Hensman et al.,, 2018), as summarized in the subsequent table (For Matérn-$\frac{3}{2}/\frac{5}{2}$, see Table 7/10). The columns of the tables indicate the VFFs for input $\tau^{\prime}$ located inside/outside $[a,b]$. The cross-covariance can be derived by substituting $h(z,\tau^{\prime})$ in (29). Table 5 and Table 6 collect the VFRFs for Matérn-$\frac{1}{2}$ LFMs with cosine/sine projection basis functions at different locations of $t^{\prime}$ in (29).
@@ D.1.2 Sine Features ($\phi(t;z)=\sin(z(t-a))$)
The VFRFs of the LFM with a Matérn-$\frac{1}{2}$ latent force are summarised in Table 5 and 6, where the absolute distances to the interval ends are denoted as $r_{a}=|t^{\prime}-a|$ and $r_{b}=|t^{\prime}-b|$ and the phase shift is $\theta=-\arctan(\frac{z}{\gamma})$. The features are continuous at $\gamma=\lambda$ when $t^{\prime}>b$.
@@
D.2 Matérn-$\frac{3}{2}/\frac{5}{2}$ Cross-covariance
Based on Table 7 and 10, we give the VFRFs for Matérn-$\frac{3}{2}$ and Matérn-$\frac{5}{2}$ LFMs with $\theta=-\arctan(\frac{z}{\gamma})$ in Table 8,9,11 and 12. Also, the derived cross-covariance expressions can return to the scaled VFFs of the latent force.
@@ D.2 Matérn-$\frac{3}{2}/\frac{5}{2}$ Cross-covariance
Appendix E Experimental Details
All models in the experimental section are implemented using GPyTorch (Gardner et al.,, 2018), trained by Adam Optimizer on an NVIDIA A100-SXM4 GPU (for TIMIT and UCI datasets) or Apple Macbook CPUs (for illustrative examples), with a learning rate of 0.01 and a batch size of 10,000. The models using doubly stochastic variational inference, e.g., IDDGPs, DLFM-VFRF, employ five samples for layer-wise sampling during training. We follow Salimbeni and Deisenroth, (2017) to set up a linear mean function for all the inner layers and a zero-mean function for the outer layer to avoid pathological behaviour (Duvenaud et al.,, 2014). The weights of the linear mean function are fixed and determined by SVD if the input and output dimensions are not equal. The variational distributions over inducing variables are initialised to normal distributions with zero mean and variances identity for the outer layers and $10^{-5}$ for the inner layers. The inducing points are initialised with K-means. All models, including RFF-based models, used 100 Monte Carlo samples for evaluations on test data.